### v2

@@ 3 Operator-relative regularity
To facilitate treating primal-dual methods as forward-backward methods with respect to suitable operators in the next section, we introduce operator-relative smoothness and monotonicity concepts. That is, we will not use the standard norms of the space, but (semi-)norms of the form $\|x\|_{\Lambda}:=\sqrt{\langle x|\Lambda x\rangle_{X,X^{*}}}$ for a positive semi-definite self-adjoint $\Lambda\in\mathbb{L}(X;X^{*})$. This will be useful for the treatment of primal-dual methods as forward-backward methods with respect to appropriate (semi-)norms. We start with the relevant definitions in Section 3.1, and then prove the relevant operator-relative descent inequalities and three-point monotonicity in Section 3.2.
@@ 3.1 Definitions
$\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle\geq q_{\Gamma}(z-\bar{x})\quad\quad\text{for}\quad q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}\quad(z\in\Omega).$
We do not at this stage assume $\Gamma$ to be positive semi-definite. We, however, call $\Gamma$ Young if there exists a self-adjoint positive semi-definite $|\Gamma|\in\mathbb{L}(X;X^{*})$ such that
@@ 3.2 Estimates
The next result provides an operator-relative generalisation of the three-point descent inequality (20). Compared to the treatment of the latter in [31, Appendix B], it is important that $x$ ($=x^{k+1}$ in the application to forward steps at $x^{k}$) is not a priori restricted to the neighbourhood $\Omega$ of $\Gamma$-monotonicity at $\bar{x}$.
@@
Proof 3.5.
Similarly to the proof of the descent inequality in Lemma 3.2, the mean value theorem applied to $\varphi(t):=F(\bar{x}+t(z-\bar{x}))$, followed by the assumed local $\Gamma$-monotonicity of $DF$, and the Young property of $\Gamma$, establishes
@@
Applying Lemma 3.2 and summing this inequality with the descent inequality it provides, we obtain the claim.
Finally, we provide a three-point monotonicity inequality that can be used instead of Lemma 3.4 when function values are not desired. The proof we present here slightly improves [7, Lemma 15.1] even in the Hilbert space scalar factor case.
@@
Arguing for $q_{\tilde{\Gamma}}$ as in the proof of Lemma 3.4, we obtain the claim.
4 Nonconvex forward-backward type methods with inexact updates
We now need to prove the convergence of outer methods for the outer/overall problem (2), given estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ by inner and adjoint methods, the latter two satisfying the tracking theory of Section 2. In this section, we do this through a convergence theory for general inexact forward backward-type methods in a normed space $X$. Our treatment encompasses primal-dual methods, seen as forward-backward methods with respect to appropriate operator-relative (semi-)norms, discussed in the previous section. We introduce such methods in Section 4.1. Then in Section 4.2 we introduce abstract growth conditions, which we will use in Sections 4.4, 4.5, 4.6 and 4.7 to prove various forms of convergence. Before that, in Section 4.3, we will verify the growth inequalities for forward-backward and primal-dual algorithms that use the tracking theory of Section 2 for (single-loop) updates of an inner problem.
4.1 General inexact forward-backward type methods
For proper $F,G:X\to\overline{\mathbb{R}}$, consider the problem
$\min_{x\in X}F(x)+G(x).$
In this subsection, and in the examples of Section 4.3, $G$ will be convex and lower semicontinuous, and $F$ Fréchet differentiable, but the general theory of Sections 4.2, 4.4, 4.5, 4.6 and 4.7 will make no such assumption.
For an initial $x^{0}$, if $X$ is Hilbert, the iterates $\{x^{k}\}_{k=1}^{\infty}$ of the basic inexact forward-backward method are generated for some step length parameter $\tau>0$ and an estimate $\widetilde{\nabla F}(x^{k})$ of $\nabla F(x^{k})$ (not necessarily the one from Section 2) by
In implicit form the method reads
$-\tau^{-1}(x^{k+1}-x^{k})\in\widetilde{\nabla F}(x^{k})+\partial G(x^{k+1}).$
(27) $0\in H(x):=F^{\prime}(x)+\partial G(x)+\Xi x$
with the implicit method
@@ 4.1 General inexact forward-backward type methods
Algorithms of the form (28) with an exact inclusion for ~$\tilde{\partial}_{k+1}$, cover many common splitting algorithms, such as Douglas–Rachford splitting (DRS) and the primal-dual proximal splitting (PDPS) of [4]; see [7, 31]. As we will see in the following examples, with an inexact inclusion, besides the inexact gradients of Section 2, the approach also covers inexact proximal maps and mismatched adjoints [17] in primal-dual methods. In exact proximal maps were used, e.g., in [33] for point source localisation in measure spaces.
Example 4.1 (Forward-backward splitting).
For forward-backward splitting with inexact $\widetilde{F^{\prime}}(x^{k})\approx F^{\prime}(x^{k})$, we take ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})$ with $M=\tau^{-1}\operatorname{Id}$ and $\Xi=0$.
Example 4.2 (Primal-dual proximal splitting).
(29) $\min_{z\in Z}\leavevmode\nobreak\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\leavevmode\nobreak\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y).$
If $f$ is convex, subject to the standard condition on the existence of $x_{0}\in\operatorname{int}\operatorname{dom}[h\circ K]\cap\operatorname{dom}[f+g]\neq\emptyset$ with $Kx_{0}\in\operatorname{int}\operatorname{dom}h$, [^1] Several relaxations are possible, include using the relative interior, or the formulas of [2]. the Fenchel–Rockafellar theorem [7, Theorem 5.11] gives rise to the necessary and sufficient first-order primal-dual optimality conditions
where $F(z,y)=f(z)$, $G(z,y)=g(z)+h_{*}(y)$, and $\Xi=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$. If $f$ is nonconvex, the necessity can be shown through, e.g., Mordukhovich subdifferentials, and their compatibility with both convex subdifferentials and Fréchet derivatives; see, e.g., [7].
Pick step length parameters $\tau,\sigma>0$. With inexact gradients for $f$, the PDPS in Hilbert spaces then reads
When $f=j\circ S_{u}$ for $S_{u}$ a PDE solution operator, and we compute $\widetilde{\nabla f}$ following Examples 2.5 and 2.7, (30) becomes the algorithm presented in [16].
To extend (30) to general normed spaces, we write it in $X=Z\times Y$ in implicit form as (28) with ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$, where
In normed spaces, we extend the above step length assumptions as follows. In the standard Hilbert setting with $M_{y}=\operatorname{Id}$ and $M_{z}=\operatorname{Id}$, we can take $K_{z}=K$ and $K_{y}=\operatorname{Id}$.
Assumption 4.2 (PDPS step length condition).
In the setting of Example 4.2, $M_{z}$ is positive semi-definite, and $K=K_{y}K_{z}$ for some $K_{z}\in\mathbb{L}(Z;V)$, $K_{y}\in\mathbb{L}(V;Y^{*})$, and a normed space $V$. Given $\lambda\geq 0$, the step length parameters $\tau,\sigma>0$ satisfy
Lemma 4.3 (PDPS preconditioning operator).
If Section 4.1 holds, then $M$ is positive semi-definite and for any $\gamma_{z},\gamma_{y}\geq 0$ and $\gamma:=\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}/2$, we have
Proof 4.4.
By a simple application of Young’s inequality and Section 4.1, we have
for any $x=(z,y)\in Z\times Y$. This establishes the first claimed inequality. The second follows by using Young’s inequality and Section 4.1 to establish
Remark 4.5 (Testing operators and accelerated methods).
In [31, 7], “testing operators” $Z_{k}\in\mathbb{L}(X^{*};X^{*})$ are used to encode convergence rates, and to prove “accelerated” $O(1/N^{2})$ rates for the PDPS under mere primal strong convexity. They could be incorporated into our treatise, however, for simplicity, we have chosen not to do this.
4.2 Inexact growth inequalities
We now make precise the approximate inclusion in (28). For the relevant definitions, we first define the Lagrangian gap functional
Example 4.6.
Example 4.7.
For the PDPS of Example 4.2, with $x=(y,z)$, we expand
Assumption 4.7.
  1. (i)
    (32) ~$\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{\mathrm{desc},k}.$
  2. (ii)
  3. (iii)
  4. (iv)
    For some $\tilde{\eta}<\eta$, we have
Remark 4.8.
If $\Omega=X$, convergence will be global. In the examples of Section 2.1, $\Omega\neq X$ may arise from $S_{u}$, $G$, or $J$ being only locally Lipschitz continuously differentiable.
Example 4.9 (Descent inequality and gap sum for basic forward-backward splitting).
[…]
We will also need the approximations ~$\tilde{\partial}_{k+1}$ to become better as the distance between the iterates shrinks, in the sense of
Assumption 4.9.
For $H$ defined in (27), we have
(33) ~$\langle\tilde{\partial}_{k+1}-H(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}),$
for all $k\in\mathbb{N}$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ for an open neighbourhood $\Omega_{\bar{x}}$ of $\bar{x}$, a positive semi-definite self-adjoint $\breve{\Lambda}\in\mathbb{L}(X;X^{*})$, errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$, and a $\gamma\geq 0$.
(34) ~$\langle\tilde{\partial}_{k+1}|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
[…]
The errors will also need to have a finite sum:
Assumption 4.10.
  1. (a)
    (33) holds, and $\bar{x}\in H^{-1}(0)$; or
  2. (b)
    (34) holds, and $\inf_{x\in\Omega_{\bar{x}}}\mathcal{G}(x;\bar{x})\geq 0.$
[…]
4.3 Specific cases
We now explicitly verify Sections 4.2, 4.2 and 4.2 for both basic forward-backward splitting and the PDPS, as well as their inexact versions based on the estimation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ formed using inner and adjoint algorithms satisfying the tracking theory of Section 2.
We start with an abstract result for “operator-relative“ inexact forward-backward. It will form the basis of verification for both forward-backward and primal-dual methods. We will take
(36) $d_{X}(x,\tilde{x})=\|x-\tilde{x}\|_{M},\enskip d_{X^{*}}(x^{*},\tilde{x}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*}-\tilde{x}^{*})]^{1/2},\enskip b_{X}(x,\tilde{x})=\|x-\tilde{x}\|_{\Lambda}.$
Then the Fenchel–Young inequality and homogeneity ensure that $d_{X}^{*}$ is Young to $d_{X}$, as defined in Section 2.3. If $M$ is invertible, this gives $d_{X^{*}}(x^{*},\tilde{x}^{*})=\|x^{*}-\tilde{x}^{*}\|_{M^{-1}}$.
Theorem 4.11.
On a normed space $X$, for a self-adjoint and positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$, suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. For all $k\in\mathbb{N}$, construct $\widetilde{F^{\prime}}(x^{k})$ obeying Section 2.1 for the distances (36) and an $\Omega\subset X$. In (28), ensure ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$. Let $\varsigma_{p}$ be as defined in (14), $\kappa$ in (11), and $r_{\mathrm{desc}}$ and $r_{p}$ in Sections 4.2 and 4.2. Then:
  1. (i)
    Section 4.2 holds for any $\tilde{\gamma}>0$, $\eta>\tilde{\eta}\geq 0$, $p\in[1,\kappa)$, and $\varepsilon_{\mathrm{desc},k}=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$ provided $\Xi=0$, $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, $\inf[F+G]>-\infty$, and
@@ 4.3 Specific cases
    Section 4.2 holds if $\Lambda\leq cM$ for a $c>0$.
Suppose further that $G$ is $\Gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\Gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}_{M}(\bar{x},\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\Gamma_{F},\Gamma_{G}\in\mathbb{L}(X;X^{*})$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. Set $\varepsilon_{k}(\bar{x})=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$ and define $r_{p}$ by (35). If $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ with $r_{p}<\delta^{2}/2$, then, for any $\eta\geq 0$:
@@
    Section 4.2 option Item (a) holds if, for $\tilde{\Gamma}_{F}:=\Gamma_{F}-(\zeta/2)\Lambda$ and $\zeta,\beta>0$,
@@
    Section 4.2 option Item (b) holds if $\Omega$ is convex, and, for $\beta>0$,
@@
Proof 4.12.
Item (i): By Lemma 3.2, we have
@@
Combining this with Theorem 2.14 for $\bar{x}=x^{k}$ and $x=x^{k+1}$ establishes (compare (19))
$\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}(x^{k+1})$
with $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})<\infty$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$. Further combining with the subdifferentiability of $G$, we verify (32). Since we assume $\breve{\Lambda}\leq 2(1-\eta)M$ and take $\varepsilon_{\mathrm{desc},k}\propto e_{p,k}(x^{k+1})$, this verifies Section 4.2Items (i) and (ii). Because $\Xi=0$, Item (iii) requires $[F+G](x^{N})\leq r_{\mathrm{desc}}+[F+G](x^{0})$ to imply $x^{N}\in\Omega$. This holds whenever $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, as we have assumed. Likewise, we prove Item (iv) with the lower bound $\inf[F+G]-[F+G](x^{0})>-\infty$.
Item (ii): Theorem 2.16 with $x^{*}=F^{\prime}(x^{k+1})$ and $\vartheta=1$ establishes for any $C>0$ that
(37) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}\leq C\implies\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty$
and, since $F$ is $\Lambda$-firmly Lipschitz, together with (24) and $\Lambda\leq cM$ it establishes that
(38) $\frac{1}{2}\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}^{2}\leq c\|\Lambda\|\|x^{k+1}-x^{k}\|_{M}^{2}+e_{\mathrm{lip},k}.$
The antecedent of (37) implies $\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}\to 0$ via (38). Since $F$ is $\Lambda$-firmly Lipschitz, taking
$x_{k+1}^{*}=F^{\prime}(x^{k+1})-F^{\prime}(x^{k})+\tilde{\partial}_{k+1}\in F^{\prime}(x^{k+1})+\partial G(x^{k+1})+\Xi x^{k+1}=H(x^{k+1}),$
we estimate
$\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}^{2}\leq\|F^{\prime}(x^{k+1})-F^{\prime}(x^{k})\|_{X^{*}}^{2}.$
Thus the antecedent of Section 4.2 implies its consequent, verifying the assumption.
For the verification of both Items 3 and 4, we observe that (35) holds because the respective lower bounds $\gamma\geq(p-1)/2$ and $\gamma\geq p-1$ guarantee $p\leq\overline{p}_{\gamma}$, and we have explicitly assumed $1\leq p$ and $\delta^{2}>2r_{p}$, where, by our choice of $\varepsilon_{k}(\bar{x})=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$, the definition of $r_{p}$ in Section 4.2, and Theorem 2.14 and
$r_{p}:=\sup_{N\in\mathbb{N}}p^{-N}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})/(2\tilde{\gamma})\leq\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})/(2\tilde{\gamma})<\infty.$
We have also explicitly assumed the remaining neighbourhood conditions of Section 4.2, as well as $0\leq\breve{\Lambda}\leq(1-\eta)M$, so only need to verify the respective (33) or (34).
Item 3: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma 3.6 and our assumption $\breve{\Lambda}\geq 0$, since $x^{k}\subset\Omega_{\bar{x}}$, we have
@@
Combining this with Theorem 2.14 and the definition of $\breve{\Lambda}$ gives
@@
We verify (33) by combining this with the $\Gamma_{G}$-monotonicity of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|$.
Item 4: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma 3.4, since $x^{k}\subset\Omega_{\bar{x}}$, we have
@@
Combining this with Theorem 2.14 gives and the definition of $\breve{\Lambda}$ gives (compare (21))

### v3

@@ 5 Operator-relative regularity
To facilitate treating primal-dual methods as forward-backward methods with respect to suitable operators in the next section, we introduce operator-relative smoothness and monotonicity concepts. That is, we will not use the standard norms of the space, but (semi-)norms of the form $\|x\|_{\Lambda}:=\sqrt{\langle x|\Lambda x\rangle_{X,X^{*}}}$ for a positive semi-definite self-adjoint $\Lambda\in\mathbb{L}(X;X^{*})$. This will be useful for the treatment of primal-dual methods as forward-backward methods with respect to appropriate (semi-)norms. We start with the relevant definitions in Section˜5.1, and then prove the relevant operator-relative descent inequalities and three-point monotonicity in Section˜5.2.
@@ 5.1 Definitions
$\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle\geq q_{\Gamma}(z-\bar{x})\quad(z\in\Omega).$
where we write
(46) $q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}$
when $\|\,\boldsymbol{\cdot}\,\|_{\Gamma}$ may not be a well-defined (semi-)norm. We do not at this stage assume $\Gamma$ to be positive semi-definite. We, however, call $\Gamma$ Young if there exists a self-adjoint positive semi-definite $|\Gamma|\in\mathbb{L}(X;X^{*})$ such that
@@ 5.2 Estimates
The next result provides an operator-relative generalisation of the three-point descent inequality (10). Compared to the treatment of the latter in [32, Appendix B], it is important that $x$ ($x^{k+1}$ in the application to forward steps at $x^{k}$) is not a priori restricted to the neighbourhood $\Omega$ of $\Gamma$-monotonicity at $\bar{x}$.
@@
Proof 5.5.
Similarly to the proof of the descent inequality in Lemma˜5.2, the mean value theorem applied to $\varphi(t):=F(\bar{x}+t(z-\bar{x}))$, followed by the assumed local $\Gamma$-monotonicity of $DF$, and the Young property of $\Gamma$, establishes
@@
Applying Lemma˜5.2 and summing this inequality with the descent inequality it provides, we obtain the claim.
Finally, we provide a three-point monotonicity inequality that can be used instead of Lemma˜5.4 when function values are not desired. The proof we present here slightly improves [7, Lemma 15.1] even in the Hilbert space scalar factor case.
@@
Arguing for $q_{\tilde{\Gamma}}$ as in the proof of Lemma˜5.4, we obtain the claim.
6 Outer algorithms
We now explicitly verify Sections˜4.2, 4.2 and 4.2 for both basic forward-backward splitting and the PDPS, as well as their inexact versions based on the estimation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ formed using inner and adjoint algorithms satisfying the tracking theory of Section˜2. We first provide in Section˜6.1 a general result for opeartor-relative forward-backward type methods for (21). This forms the basis of verification of the assumptions both for forward-backward splitting in Section˜6.2, and primal-dual methods in Section˜6.3. In the final Section˜6.4 we also treat adjoint mismatch as another example of our general approach to inexactness.
6.1 A general result
Let $M,\Lambda\in\mathbb{L}(X;X^{*})$ be self-adjoint postive and semi-definite, with $M$ preconditioning operator from Section˜4. To use the tracking theory of Section˜2, we will impose the choices
(48) ${d_{X}(x,\tilde{x})}=\|x-\tilde{x}\|_{M},\enskip{d_{X^{*}}(x^{*},\tilde{x}^{*})}=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*}-\tilde{x}^{*})]^{1/2},\enskip{b_{X}(x^{k+1},x^{k})}=\|x^{k+1}-x^{k}\|_{\Lambda}.$
Then the Fenchel–Young inequality and homogeneity ensure that $d_{X^{*}}$} is Young to $d_{X}$}, as defined in Section˜2.2. If $M$ is invertible, ${d_{X^{*}}(x^{*},\tilde{x}^{*})}=\|x^{*}-\tilde{x}^{*}\|_{M^{-1}}$. More generally, we have the following bound.
Lemma 6.1.
With $d_{X^{*}}$ as in (48), $\|x^{*}-\tilde{x}^{*}\|_{X^{*}}\leq\|M\|_{\mathbb{L}(X;X^{*})}^{1/2}{d_{X^{*}}(x^{*},\tilde{x}^{*})}$ for all $x^{*},\tilde{x}^{*}\in X^{*}$.
Proof 6.2.
We have $f^{*}\leq g^{*}$ for $f\geq g$ from the definition of the Fenchel conjugate. Abbreviating $m:=\|M\|_{\mathbb{L}(X;X^{*})}$, and using [7, Lemma 5.8 and Lemma 5.5], it follows for any $x^{*}\in X^{*}$ that
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*})\geq(\tfrac{m}{2}\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(x^{*})=m(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(x^{*}/m)=\frac{m}{2}\|x^{*}/m\|_{X^{*}}^{2},=\frac{1}{2m}\|x^{*}\|_{X^{*}}^{2}.$
[…]
In brief, the next theorem says that
  1. 1.
    Sections˜4.2 and 4.2, used for subdifferential convergence by Theorem˜4.6, require that the initialisation $x^{0}$ be in the set $\Omega$ where the tracking assumptions hold, and that the step lengths (encoded in $M$) be small compared to the operator-relative Lipschitz factor $\Lambda$.
  2. 2.
    Section˜4.2, used for stronger convergence results by Corollaries˜4.12, 4.13 and 4.16, also requires sufficient strong subdifferentiability.
Theorem 6.3.
On a normed space $X$, let $M,\Lambda\in\mathbb{L}(X;X^{*})$ be self-adjoint and positive semi-definite. Suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Given an initial $x^{0}\in X$, for all $k\in\mathbb{N}$, construct $\widetilde{F^{\prime}}(x^{k})$ obeying Section˜2.1 (or, more generally, Theorems˜2.5 and 2.8 without Section˜2.1) for the distances (48) and an $\Omega\subset X$. Update $x^{k+1}$ by solving
(49) $0\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}+M(x^{k+1}-x^{k}).$
Let $\varsigma_{p}$, $\kappa$ and $e_{p,k}$ be as in Theorem˜2.5. Then,
  1. (i)
@@ 6.1 A general result
    and, for some $\tilde{\gamma}>0$,
@@
    Section˜4.2 holds if $\Lambda\leq cM$ for a $c>0$.
Suppose further that $G$ is $\Gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\Gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}_{M}(\bar{x},\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\Gamma_{F},\Gamma_{G}\in\mathbb{L}(X;X^{*})$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If
$x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})\quad\text{with}\quad r_{p}=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}e_{p,k}/(2\tilde{\gamma})<\delta^{2}/2,$
[…]
@@
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\Gamma}_{F}:=\Gamma_{F}-(\zeta/2)\Lambda$ and some $\zeta,\beta>0$,
@@
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$,
@@
Proof 6.4.
Item˜(i): By Lemma˜5.2, we have
@@
Combining this with Theorem˜2.5 for $\bar{x}=x^{k}$ establishes (compare (9))
$\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}$
with $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}<\infty$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$. Further combining with the subdifferentiability of $G$, we verify (26). Since we assume $\breve{\Lambda}\leq 2(1-\eta)M$ and take $\varepsilon_{\mathrm{desc},k}\propto e_{p,k}$, this verifies Section˜4.2Items˜(i) and (ii). Because $\Xi=0$, Item˜(iii) requires $F+G](x^{N})\leq r_{\mathrm{desc}}+[F+G](x^{0})$ to imply $x^{N}\in\Omega$. This holds whenever $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, as we have assumed. Likewise, we prove Item˜(iv) with the lower bound $\inf[F+G]-[F+G](x^{0})>$-$\infty$.
Item˜(ii): We have $\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}\leq\|F^{\prime}(x^{k+1})-\widetilde{F^{\prime}}(x^{k})\|_{X^{*}}$ through the choice
$x_{k+1}^{*}=F^{\prime}(x^{k+1})-\widetilde{F^{\prime}}(x^{k})+\tilde{\partial}_{k+1}\in F^{\prime}(x^{k+1})+\partial G(x^{k+1})+\Xi x^{k+1}=H(x^{k+1}).$
Hence, the triangle inequality, Eqs.˜45 and 6.1 followed by Theorem˜2.8 establish
(50) $\begin{split}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}&\leq\|F^{\prime}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}+\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})\|_{X^{*}}\\[-8.61108pt] &\leq\|\Lambda\|^{1/2}\|x^{k+1}-x^{k}\|_{\Lambda}+\|M\|_{\mathbb{L}(X;X^{*})}^{1/2}{d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\\ &\leq\|\Lambda\|^{1/2}\|x^{k+1}-x^{k}\|_{\Lambda}+\left(\|M\|_{\mathbb{L}(X;X^{*})}e_{\mathrm{lip},k}\right)^{1/2},\end{split}$
where $e_{\mathrm{lip},k}\geq 0$ satisfies
(51) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}<\infty\implies\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty.$
The antecedent of (51) implies $\lim_{k\to\infty}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}=0$ through Eqs.˜50 and 51. Since $\Lambda\leq cM$ by assumption, this establishes Section˜4.2.
[…]
We have also explicitly assumed the remaining neighbourhood conditions of Section˜4.2, as well as $0\leq\breve{\Lambda}\leq(1-\eta)M$, so only need to verify the respective (27) or (28).
Item˜3: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma˜5.6 and our assumption $\breve{\Lambda}\geq 0$, since $x^{k}\subset\Omega_{\bar{x}}$, we have
@@
Here we recall from (46) the notation $q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}$, used when $\|\,\boldsymbol{\cdot}\,\|_{\Gamma}$ may not be a well-defined (semi-)norm. Combining this with Theorem˜2.5 and the definition of $\breve{\Lambda}$ gives
@@
Since $G$ is $\Gamma_{G}$-monotone and $\Xi$ is skew-symmetric, we have
$\langle\partial G(x^{k+1})-\partial G(\bar{x})+\Xi(x^{k+1}-\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\|x^{k+1}-\bar{x}\|_{\Gamma_{G}}.$
Due to (49), we have $M(x^{k+1}-x^{k})=\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$. By definition, $H(x)=\partial G(x)+F^{\prime}(x)+\Xi x$. Hence, summing the previous two inequalities, it follows that
$\langle\tilde{\partial}_{k+1}-H(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\Gamma_{G}+\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
After an application of the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|$, this verifies (27) through the.
Item˜4: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma˜5.4, since $x^{k}\subset\Omega_{\bar{x}}$, we have
@@
Combining this with Theorem˜2.5 gives and the definition of $\breve{\Lambda}$ gives (compare (11))